# Error correction of transversal controlled-NOT gates for scalable surface code computation

### v1

@@ V.2 A comparison with lattice surgery
In previous sections, we have laid out our scalable tCNOT strategy. In this section, we turn to an examination of lattice surgery [horsman_surface_2012, fowler_low_2018, litinski_lattice_2018, litinski_game_2019, chamberland_universal_2022]. In lattice surgery, static logical surface code patches are set up with bridging regions of unentangled ancilla qubits between them. Stabilizer measurements on these bridging regions are turned on and off to connect the logical operators of individual surface codes and perform joint logical Pauli measurements, where reliable measurement of the joint Pauli operators requires $d$ rounds of stabilizer measurements. Arbitrary logical Clifford gates can be executed via combinations of these joint measurements and logical Pauli gates.
The decoding of these operations has been well studied, being nearly identical to decoding a SCQM. As in the case of a SCQM, a conventional MWPM decoder works well, and thus we use this decoder for comparison. We refer the reader to Appendix  D-F for details. In practice, a lattice-surgery based quantum algorithm is compiled into the shortest sequence of joint measurements instead of directly using a gate set comprising of CNOT gates [litinski_game_2019, fowler_optimal_2013], and a transversal implementation may use logical blocks with $m>1$. In this context, we leave the somewhat artificial comparison of an isolated fault-tolerant tCNOT versus a lattice-surgery CNOT to App. G, displaying the numerical results in Fig. 4 (center). Here, we compare the cost of a state teleportation circuit implemented using both approaches that may be directly used in a logical quantum algorithm.
In Fig. V(b), we present the joint-parity version of the teleportation circuit in Fig. V(a). Since only a $ZZ$ logical measurement between the two surface codes is required, this operation takes $d$ rounds and requires no additional logical ancilla patches, which matches the overhead of an isolated fault-tolerant tCNOT within a logical quantum algorithm. This suggests that in certain settings, the two strategies may use equivalent resources. We leave an extended overhead analysis to future work, focusing on fault-tolerant thresholds for both approaches in the next section.
@@ VI Logical gates for erasure qubits
Until now, we have analysed logical gates under i.i.d. Pauli noise. We now move onto a corresponding analysis for qubits where the dominant errors include a form of structured noise called erasures [grassl_codes_1997, wu_erasure_2022]. Erasures consist of detectable errors that result in the affected qubit being projected into the maximally mixed state. We further investigate a more tailored model consisting of biased erasures [sahay_high-threshold_2023], where detectable errors only happen from one half of the computational subspace. We review the biased erasure noise model in App. H.
Figure 7: MWPM thresholds under circuit-level noise while varying the erasure fraction $R_{e}$ for two logical operations: a lattice-surgery CNOT (browns), and a tCNOT with ordered decoding (blues).
@@
We analyse the change in MWPM-based thresholds with $R_{e}$ for logical operations in Fig. 7. Fig. 7(top) shows results for conventional erasures, and Fig. 7(bottom) for biased erasures. We briefly summarize the results for biased erasures in the following: the threshold for a tCNOT corrected using ordered decoding increases from $1.03\%$ at $R_{e}=0$ to $8.3\%$ at $R_{e}\approx 1$. The thresholds for joint-measurements (denoted by LS-CNOTs in Fig. 7) exactly coincide with a SCQM up to error bars, increasing from $1.04\%$ at $R_{e}=0$ to $10.3\%$ at $R_{e}\approx 1$. These values also match with those of ordered decoding on a tCNOT except at extremely high erasure fractions $R_{e}>0.95$. Concretely, at $R_{e}=0.98$ we see that $p_{t}=7.5\%$ for ordered decoding, while $p_{t}=8.3\%$ for regular decoding used during lattice surgery. A brief explanation for this phenomenon is given in Appendix C.1. We note that the fact that tCNOT thresholds for dominant erasure noise are strictly lower than the corresponding joint-measurements thresholds will ultimately affect the relative error rates of transversal logical operations.
@@ VII Conclusion
In this work, we have performed an analysis of error correction of transversal CNOTs (tCNOTs) on surface codes in the context of scalable quantum computation. We highlight the necessity for gadget fault-tolerance in large-scale quantum algorithms. In this context, we present various decoding strategies that may be used to correct these operations, highlighting an intuitive strategy, ordered decoding, that uses the deterministic propagation of Pauli errors through the logical gate. This strategy is compared with previous proposals to correct transversal logical operations - combined hypergraph decoding, and single-update decoding, showing that ordered decoding maintains the graph-based error correction advantages and thresholds of surface code memory experiments.
@@
We now briefly turn to a discussion focusing on hardware limitations. Most metrics we study indicate that where permitted by the architecture, a transversal strategy is more advantageous: tCNOTs corrected using ordered decoding have similar thresholds and lower logical error rates compared to lattice surgery-based approaches. There are, however, some subtle caveats. Transversal CNOT gates intrinsically involve nonlocal connectivity. Realistically, long-range interactions for a particular hardware may be slower and/or exhibit higher error rates than nearest-neighbour interactions. Specifically for the example of reconfigurable neutral atoms, the corresponding error contribution arises from qubit movement across distances scaling with the surface code size [bluvstein_quantum_2022, evered_high-fidelity_2023]. As another example, for static neutral atom systems with long-range interactions realised via the Rydberg blockade, two-qubit gate fidelity decays quasi-linearly with the gate range [pecorari_high-rate_2024, poole_architecture_2024], setting a maximum radius over which using such non-local connectivity is practical, beyond which a threshold disappears. Thus, the relative performance of these gate strategies will vary significantly based on hardware constraints.
Our work thus builds towards an analysis of fault-tolerant transversal logical operations for use in quantum algorithms to be implemented in hardware. Further study of transversal non-Clifford gates [moussa_transversal_2016, brown_fault-tolerant_2020], decoders for logical gadgets, and optimal compilation schemes [fowler_optimal_2013] will indicate the performance of these strategies as a whole.
@@ VIII Acknowledgements
We are grateful to Yue Wu, Shraddha Singh, Pei-Kai Tsai, and Aleksander Kubica for helpful discussions. We particularly thank Jahan Claes for insightful contributions during early stages of this project. We acknowledge the Yale Center for Research Computing for use of the Grace cluster. This work was supported by the National Science Foundation (QLCI grant OMA-2120757). Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the authors and do not necessarily reflect the views of NSF. After the completion of this work, we became aware of a related decoder implementation investigating low-depth circuits that uses the same underlying principle as ordered decoding [wan_iterative_2024].
Appendix A Hyperedges in the tCNOT
@@
With this set of detectors, it is possible to extract independent error mechanisms that create three defects. For example, consider a measurement error on $T$ at round $d$. Assuming no other errors, $S_{TX}^{d}=1$ and all other outcomes are 00, meaning that the set of detectors $\{v_{TX}^{{d}},v_{CX}^{{d+1}},v_{TX}^{{d+1}}\}=1$ produce defects. This 3-component term is a hyperedge.
Appendix B Graph decoders for the tCNOT
@@
/* Updating dependent measurement outcomes */
@@ 2       ${S}^{i}_{dep}\leftarrow{S}^{i}_{dep}\oplus{S}^{i}_{ind}$
/* Updating dependent defect nodes */
4 foreach ${v}_{S}^{i}\in G_{dep}$ do
@@ 10              $p_{e}\;\leftarrow p_{e}+p_{e_{G_{ind}}}-2p_{e}p_{e_{G_{ind}}}$
We note a fundamental conceptual similarity between ordered decoding and the two-pass or iterative MWPM decoding methods [fowler_optimal_2013, yuan_modified_2022, paler_pipelined_2023, iolius_performance_2023]. These approaches have been used to improve SCQM decoding accuracy by leveraging probabilistic intra-SC correlations. In contrast, for ordered decoding, we make use of deterministic inter-SC error correlations induced by the tCNOT.
@@ 1 Decode $G_{ind}$ independently, obtaining the set of matched defect pairs $\mathcal{E}^{G_{ind}}$
2 foreach unordered pair $\mathcal{E}=(v_{S_{k}}^{i},v_{S_{l}}^{j})\in\mathcal{E}^{G_{ind}}$ do
3       if $\exists v_{S_{k}}^{i}\in\mathcal{E}\mid i\leq d$ then
@@
Decode $G_{dep}$ independently
@@
Appendix C Uncorrectable propagated errors for ordered decoding on the tCNOT
Here, we discuss the logical error rates observed for ordered decoding using a matching subroutine. For $Z$ errors, ordered decoding uses the simple concept of a target-first-control-next decoding strategy. The first round of decoding on $T$ pairs up defects on the $X$ stabilizer graph. These defect pairs correspond to errors that may or may not be copied over to $C$ by the tCNOT. Assuming errors are corrected up to stabilizers, these defect pairs arising from underlying physical errors and their corresponding predicted errors returned by a decoder create space-time cycles in the decoding graph. These cycles can be broadly classified into three distinct categories:
@@ Appendix C Uncorrectable propagated errors for ordered decoding on the tCNOT
    Spacetime cycles spanning the tCNOT round: This defect pattern arises from a combination of measurement and data errors. There are multiple minimal-weight ways of projecting such a path. For example, see Fig. 8. Suppose the decoder identifies the grey path with the ‘data-qubit projection’ entirely below the tCNOT, but the actual physical error was the red error with the ‘data-qubit projection’ above the tCNOT. The former will transmit errors to the target, but the latter will not. Given the decoder’s choice, the target will be updated incorrectly. Hereon, we refer to this class of errors as ambiguous errors.
Figure 8: Errors can span the tCNOT measurement round, creating a defect below the tCNOT paired to one above it. Depending on where the data qubit errors actually occurred, different errors are copied over to $G_{dep}$ for the same defect pattern. If the error in red occurred, $C$ remains unaffected.
Ambiguous errors arising from spacetime cycles spanning the tCNOT round are effectively uncorrectable by an ordered decoder - the decoder can only randomly choose the truly correct data-error projection. As a result, ambiguous errors lead to an increase in the logical error rate of the dependent graphs. However, from Fig. 4, we observe that they do not materially degrade decoder thresholds.
C.1 Ordered decoding at high $R_{e}$
Here, we discuss the apparent reduction in threshold at high erasure fractions for a tCNOT with ordered decoding compared to an SCQM experiment. In essence, this is because of ambiguous errors that create cycles around the tCNOT round. For all other categories of correctable erasure errors not spanning the tCNOT round, the decoder can use the erasure flag on qubits suffering errors to determine the exact path of errors propagating over onto $G_{dep}$.
Observe the example Fig. 8. Unlike the previous section, we envision that erasure errors have occurred along both grey and red paths. However, only one (red) path truly has $Z$ errors due to the erasures, giving rise to two defects. Exactly like the case with Pauli errors, the decoder has no way to determine which of the two error patterns occurred. If it guesses the grey path occurred, an incorrect update is applied to the control. At high $p$ and $R_{e}$, the probabilities of such cycles increase. At sufficiently high $pR_{e}$ logical errors are dominated by these ambiguous cycles for which the decoding ability is Pauli-like instead of erasure-like. This in turn reduces $p_{t}$ for the tCNOT at high erasure fractions.
Appendix D Logical computation using lattice surgery

### v2

@@ V.2 A comparison with lattice surgery
In previous sections, we have laid out our scalable tCNOT strategy. In this section, we turn to a comparison with lattice surgery [horsman_surface_2012, fowler_low_2018, litinski_lattice_2018, litinski_game_2019, chamberland_universal_2022]. In lattice surgery, static logical surface code patches are set up with bridging regions of unentangled qubits between them. Stabilizer measurements on these bridging regions are turned on and off to connect the logical operators of individual surface codes and perform joint logical Pauli measurements. Reliable measurement of these joint Pauli operators requires $d$ rounds of stabilizer measurements. Arbitrary logical Clifford gates can be executed via combinations of these joint measurements and Pauli gates.
The decoding of these operations has been well studied, being nearly identical to decoding a SCQM. As in the case of a SCQM, a conventional MWPM decoder works well, and thus we use this decoder for comparison. We refer the reader to Appendix  E-G for details. In practice, a lattice-surgery based quantum algorithm is compiled into the shortest sequence of joint measurements instead of directly using a gate set comprising of CNOT gates [litinski_game_2019, fowler_optimal_2013], and a transversal implementation may use logical blocks with $m,g>1$. In this context, we leave the somewhat artificial comparison of an isolated fault-tolerant tCNOT versus a lattice-surgery CNOT to App. H, showing therein that the tCNOT has both lower overheads and logical error rates. We display the numerical results in Fig. 4. Here, we compare the cost of a state teleportation circuit implemented using both approaches that may be directly used in a logical quantum algorithm.
In Fig. V(b), we show the joint-parity version of the teleportation circuit in Fig. V(a). Since only a $ZZ$ logical measurement between the two surface codes is required, this operation takes $d$ rounds and requires no additional logical ancilla patches, which matches the overhead of an isolated fault-tolerant tCNOT within a logical algorithm. This suggests that in certain settings, the two strategies may use equivalent resources. We leave an extended overhead analysis to future work, focusing on fault-tolerant thresholds for both approaches in the next section.
@@ VI Logical gates for erasure qubits
Until now, we have analysed logical gates under i.i.d. Pauli noise. We now move onto a corresponding analysis for qubits where the dominant errors include a form of structured noise called erasures [grassl_codes_1997, wu_erasure_2022]. Erasures consist of detectable errors that result in the affected qubit being projected into the maximally mixed state. We further investigate a more tailored model consisting of biased erasures [sahay_high-threshold_2023], where detectable errors only happen from one half of the computational subspace. We review the biased erasure noise model in App. J.
Figure 7: MWPM thresholds under circuit-level noise while varying the erasure fraction $R_{e}$ for a lattice-surgery CNOT (browns), and a tCNOT with ordered decoding (blues).
@@
We analyse the change in MWPM-based thresholds with $R_{e}$ for logical operations in Fig. 7. Fig. 7(top) shows results for conventional erasures, and Fig. 7(bottom) for biased erasures. We briefly summarize the results for biased erasures in the following: the threshold for a tCNOT corrected using ordered decoding increases from $1.03\%$ at $R_{e}=0$ to $8.3\%$ at $R_{e}\approx 1$. The thresholds for joint-measurements (denoted by LS-CNOTs in Fig. 7) exactly coincide with a SCQM up to error bars, increasing from $1.04\%$ at $R_{e}=0$ to $10.3\%$ at $R_{e}\approx 1$. These values also match with those of ordered decoding on a tCNOT except at extremely high erasure fractions $R_{e}>0.95$. Concretely, at $R_{e}=0.98$ we see that $p_{t}=7.5\%$ for ordered decoding, while $p_{t}=8.3\%$ for regular decoding used during lattice surgery. A brief explanation for this phenomenon is given in Appendix D.1. The fact that tCNOT thresholds for dominant erasure noise are strictly lower than the corresponding joint-measurements thresholds will ultimately affect the relative error rates of erasure-dominated transversal logical operations.
@@ VII Conclusion
In this work, we have performed an analysis of error correction of transversal CNOTs (tCNOTs) on surface codes in the context of scalable quantum computation. We highlight the utility of gadget fault-tolerance in large-scale quantum algorithms. In this context, we present a unified framework to describe various decoding strategies that may be used to correct tCNOT operations, focusing on an intuitive strategy, ordered decoding, that uses the deterministic propagation of Pauli errors through the logical gate. This strategy is compared with previous proposals to correct transversal logical operations - combined hypergraph decoding, and single-update decoding, showing that ordered decoding maintains the graph-based error correction advantages and thresholds of surface code memory experiments.
@@
We now briefly turn to a discussion focusing on hardware limitations. Most metrics we study indicate that where permitted by the architecture, a transversal strategy is more advantageous: tCNOTs corrected using ordered decoding have similar thresholds and lower logical error rates compared to lattice surgery-based approaches. There are, however, some subtle caveats. Transversal CNOT gates intrinsically involve nonlocal connectivity. Realistically, long-range interactions for a particular hardware may be slower and/or exhibit higher error rates than nearest-neighbour interactions. Specifically for the example of reconfigurable neutral atoms, the corresponding error contribution arises from qubit movement across distances scaling with the surface code size [bluvstein_quantum_2022, evered_high-fidelity_2023]. As another example, for static neutral atom systems with long-range interactions realised via the Rydberg blockade, two-qubit gate fidelity decays quasi-linearly with the gate range [pecorari_high-rate_2024, poole_architecture_2024], setting a maximum radius over which using such non-local connectivity is practical. Thus, the relative performance of these gate strategies will vary significantly based on hardware constraints.
Our work thus builds towards an analysis of fault-tolerant transversal logical operations for use in quantum algorithms to be implemented in hardware. Further study of transversal non-Clifford gates [moussa_transversal_2016, brown_fault-tolerant_2020], decoders for logical gadgets, and optimal compilation schemes [fowler_optimal_2013] that use a unified circuit-decoder perspective will indicate the performance of these strategies as a whole.
@@ VIII Acknowledgements
We are grateful to Yue Wu, Shraddha Singh, Pei-Kai Tsai, Qile Su, and Aleksander Kubica for helpful discussions. In particular, we thank Jahan Claes for insightful contributions during early stages of this project. We acknowledge the Yale Center for Research Computing for use of the Grace cluster. This work was supported by the National Science Foundation (QLCI grant OMA-2120757). Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the authors and do not necessarily reflect the views of NSF. After the completion of this work, we became aware of a related decoder implementation investigating low-depth circuits that uses the same underlying principle as ordered decoding [wan_iterative_2024].
Appendix A Propagated defects in the dynamic frame
Here, we expand on the functioning of the single-update decoder. Recall that it operates in the dynamic frame. In terms of the measured static checks, we have the dynamic frame detectors
$v_{CX}^{{d}}=S_{CX}^{d-1}\oplus S_{CX}^{d},$ (A1a)
$v_{CX}^{{d+1}}=S_{CX}^{d}\oplus S_{CX}^{d+1}\oplus S_{TX}^{d+1},\text{ and}$ (A1b)
$v_{CX}^{{d+2}}=S_{CX}^{d+1}\oplus S_{TX}^{d+1}\oplus S_{TX}^{d+2}\oplus S_{TX}^{d+2}$ (A1c)
We illustrate the behaviour of the decoder using an instance of a data qubit error $Z_{T,q}$ on a qubit $q$ in $T$ that creates defects on $T$ at round $k$:
  1. 1.
    If $k\leq d$: After round $d$, $Z_{T,q}$ is copied over to the control, creating $Z_{C,q}$. Since $\left[Z_{T,q}Z_{C,q},\mathbf{S}_{CX}=S_{CX}S_{TX}\right]=0$, no defect is created in the dynamic frame after the tCNOT despite the actual error being present. This behaviour differs from the static frame.
  2. 2.
    If $k>d$: $Z_{T,q}$ is not copied over to $C$, but creates new defects on the updated $C$ due to the updated checks of the dynamic frame: $\left[Z_{T,q},\mathbf{S}_{CX}\right]\neq 0$. These errors are not truly on the control, and the resulting ‘false defects’ simply mirror defect patterns on the target. Note that these false defects appear in addition to defects created by errors on the target itself, resulting in an effective doubling of the defect rate post-tCNOT on $T$.
We observe that use of the dynamic frame prevents propagated errors from creating defects on $C$, while creating false defects from errors that did not propagate. After independently decoding and correcting both subgraphs in the dynamic frame, we use knowledge of all the corrections applied to $T$ to apply a final logical update to $C$ and restore the code states (Alg. 1).
Appendix B Hyperedges in the tCNOT
@@
With this set of detectors in the hybrid frame, it is possible to extract independent error mechanisms that create three defects. For example, consider a measurement error on $T$ at round $d$. Assuming no other errors, $S_{TX}^{d}=1$ and all other outcomes are 00, meaning that the set of detectors $\{v_{TX}^{{d}},v_{CX}^{{d+1}},v_{TX}^{{d+1}}\}=1$ produce defects. This 3-component term is a hyperedge.
Appendix C Graph decoders for the tCNOT
@@
/* Change dependent graphs to dynamic frame */
@@ 2       ${S}^{i}_{dep}\leftarrow{S}^{i}_{dep}\oplus{S}^{i}_{ind}$
4foreach ${v}_{S}^{i}\in G_{dep}$ do
@@ 10              $p_{e}\;\leftarrow p_{e}+p_{e{G_{ind}}}-2p_{e}p_{e_{G_{ind}}}$
We note a fundamental conceptual similarity between ordered decoding and the two-pass or iterative MWPM decoding methods [fowler_optimal_2013, yuan_modified_2022, paler_pipelined_2023, iolius_performance_2023]. These approaches have been used to improve SCQM decoding accuracy by leveraging probabilistic intra-SC correlations. In contrast, for ordered decoding, we make use of deterministic inter-SC error correlations induced by the tCNOT. It may be possible to combine the two approaches, along with other subroutines that make use of soft information such as belief propagation [higgott_improved_2023], to further improve the performance of ordered decoding.
@@ 1 Decode $G_{ind}$ independently, obtaining the set of matched defect pairs $\mathcal{E}^{G_{ind}}$
3 foreach unordered pair $e=(v_{S_{k}}^{i},v_{S_{l}}^{j})\in\mathcal{E}^{G_{ind}}$ do
4       if $\exists v_{S_{k}}^{i}\in{e}\mid i\leq d$ then
@@
9Decode $G_{dep}$ independently
/* Update $G_{dep}$’s logical status */
$L^{G_{dep}}\leftarrow L^{G_{dep}}\oplus L^{\prime}$
@@
C.1 A phenomenological threshold comparison
Hypergraph decoders on surface codes show an advantage over MWPM in the presence of $Y$ errors. This is because they can directly treat $Y$ errors using hyperedges. In contrast, an MWPM decoder must decompose $Y$ errors into independent $X$ and $Z$ errors and graphs. This decomposition results in the loss of initial $Y=XZ$ correlations. Here, we attempt to control for this possible $Y$ error advantage of HUF during tCNOT decoding.
We perform a comparison between HUF and ordered MWPM decoding for a tCNOT using a tailored phenomenological noise model. In this model, independent bit-flip errors on data qubits and measurement errors are applied with equal probability $p$ in each stabilizer measurement round. This removes $Y$ errors, allowing a direct comparison of the ability of the two decoders to correct correlated errors that arise solely from error propagation through tCNOT gates.
We study the same tCNOT system as in Sec. IV, with the results shown in Fig. A1. Similar to the circuit-level model, we find a higher threshold using an ordered MWPM decoder ($p_{t}=2.77\%$) compared to HUF ($p_{t}=2.47\%$). Additionally, the logical error rate of ordered MWPM is slightly lower than that of HUF, reversing the general trend for circuit level noise that includes $Y$ errors seen in Fig. 3. These findings suggest that (1) much of the logical error rate advantage of HUF indeed arises from its improved handling of $Y$ errors, and (2) an ordered MWPM decoder may outperform HUF when correcting correlated errors solely arising from tCNOT gates.
Figure A1: Thresholds and logical error rates of an HUF and ordered MWPM decoder for a tCNOT under biased phenomenological noise. Thresholds are found by finite size scaling close to $p_{t}$ (transculent lines).
Appendix D Uncorrectable propagated errors for ordered decoding on the tCNOT
Figure A2: Errors can span the tCNOT measurement round, creating a defect below the tCNOT paired to one above it. Depending on where the data qubit errors actually occurred, different errors are copied over to $G_{dep}$ for the same defect pattern. If the error in red occurred, $C$ remains unaffected.
Here, we discuss the logical error rates observed for ordered MWPM decoding. For $Z$ errors, ordered decoding uses the simple concept of a target-first-control-next decoding strategy. The first round of decoding on $T$ pairs up defects on the $X$ stabilizer graph. These defect pairs correspond to errors that may or may not be copied over to $C$ by the tCNOT. Assuming errors are corrected up to stabilizers, these defect pairs arising from underlying physical errors, and their corresponding predicted errors returned by a decoder create space-time cycles in the decoding graph (Fig. A2). These cycles can be broadly classified into three distinct categories:
@@ Appendix D Uncorrectable propagated errors for ordered decoding on the tCNOT
    Spacetime cycles spanning the tCNOT round: This defect pattern arises from a combination of measurement and data errors. There are multiple minimal-weight ways of projecting such a path. For example, see Fig. A2. Suppose the decoder identifies the grey path with the ‘data-qubit projection’ entirely below the tCNOT, but the actual physical error was the red error with the ‘data-qubit projection’ above the tCNOT. The former will transmit errors to the target, but the latter will not. Given the decoder’s choice, the target will be updated incorrectly. Hereon, we refer to this class of errors as ambiguous errors.
Ambiguous errors arising from spacetime cycles spanning the tCNOT round are effectively uncorrectable by the first stage of an ordered decoder - the decoder can only randomly choose the truly correct data-error projection. Subsequent error correction of the residual defects takes place on the dependent subgraph. As a result, ambiguous errors lead to an increase in the logical error rate of the dependent graphs. However, from Fig. 4, we observe that they do not materially degrade decoder thresholds.
D.1 Ordered decoding at high $R_{e}$
Here, we discuss the apparent reduction in threshold at high erasure fractions for a tCNOT with ordered decoding compared to an SCQM experiment. In essence, this is because of ambiguous errors that create cycles around the tCNOT round. For all other categories of correctable erasure errors, the decoder can use the erasure flag on qubits suffering errors to determine the exact path of errors propagating over onto $G_{dep}$.
Observe the example Fig. A2. Unlike the previous section, we envision that erasure errors have occurred along both grey and red paths. However, only one (red) path truly has $Z$ errors due to the erasures, giving rise to two defects. Exactly like the case with Pauli errors, the decoder has no way to determine which of the two error patterns occurred. If it guesses the grey path occurred, an incorrect update is applied to the control. At high $p$ and $R_{e}$, the probabilities of such cycles increase. At sufficiently high $pR_{e}$ logical errors are dominated by contributions from these ambiguous cycles for which the decoding ability is Pauli-like instead of erasure-like. This in turn reduces $p_{t}$ for the tCNOT at high erasure fractions.
Appendix E Logical computation using lattice surgery